OpenStudy (anonymous):

@Kreshnik an easy one for you \(\LARGE \lim_{x \rightarrow 8} \frac{16 - x}{x}\)

OpenStudy (anonymous):

oh man i screwed up >.<

OpenStudy (anonymous):

\[\Huge \lim_{x\to 8}\frac{16-x}{x}\] warm up? lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ill just post the real one next time

OpenStudy (anonymous):

nahh... it's just about substituting.. ... \[\LARGE \lim_{x\to 8}\frac{16-x}{x}=\frac{16-8}{8}=\frac88=1\] ... ok my turn? :) (level 1 too :P )

OpenStudy (anonymous):

\[\LARGE \lim_{x\to 3}\frac{x^2-9}{x+3}\]

OpenStudy (anonymous):

this can be solve two ways but to avoid all the latex ill go with 3^2 - 9/ 4+3 = 9 - 9/6 = 0/6 = 0

OpenStudy (anonymous):

that's correct ...3^2 - 9/ 4+3 i guess here yo meant 3^2 - 9/ 3+3 ;) typo... Your turn ;)

OpenStudy (anonymous):

1 question... am I allowed to use paper? (to calculate on my notebook)

OpenStudy (anonymous):

why not hehe

OpenStudy (anonymous):

??

OpenStudy (anonymous):

this is hard hmmm

OpenStudy (anonymous):

please dont... lol hahaha...

OpenStudy (anonymous):

\(\Large \lim_{h \rightarrow 0} \frac{(x+ h)^2 - 2(x+h) + 1 - x^2 + 2x - 1}{h}\) think you can handle that? :D

OpenStudy (anonymous):

multiple choices?? (I've never done such thing like that... I'll try though )

OpenStudy (anonymous):

lol it's where derivatives come from... \(\LARGE \lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{h}\) but it's easy..lemme give some choices

OpenStudy (anonymous):

is it 2x-2

OpenStudy (anonymous):

haha you got it :DDD

OpenStudy (anonymous):

\[\Large \lim_{h\to0}\frac{(x+h)^2-2(x+h)+1-x^2+2x-1}{h}\] \[\Large \lim_{h\to0}\frac{x^2+2xh+h^2-2x -2h-x^2+2x }{h}\] \[\Large \lim_{h\to0}\frac{\cancel x^2+2xh+h^2-\cancel {2x} -2h-\cancel x^2+\cancel {2x} }{h}\] \[\Large \lim_{h\to0}\frac{h(2x+h -2) }{h}\] 2x+0-2 2x-2 :)

OpenStudy (anonymous):

<pause> im gonna eat..

OpenStudy (anonymous):

ok.. boun apetit ;)

OpenStudy (anonymous):

when you're back, let me know. :)

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

ok... should I continue with a bit harder limit or still warming up ? :D

OpenStudy (anonymous):

you can go serious now :p

OpenStudy (anonymous):

\[\Huge \lim_{x\to 1}\frac{\sqrt[4]{x}-1}{\sqrt[3]{x}-1}\] A) 0 B) 2/3 C) 3/4 D) 4/3 Good Luck..

OpenStudy (anonymous):

huh..you outdid yourself this time :D

OpenStudy (anonymous):

can you handle that? :D.. need hints just tell me :D

OpenStudy (anonymous):

how do you mean, "outdid" ... can you explain it.

OpenStudy (anonymous):

would you like me to change the question? (I didn't copy this on book, it was on my mind.) next one is too... :D

OpenStudy (anonymous):

it's a good one

OpenStudy (anonymous):

but it's a long one too.. !! lol

OpenStudy (anonymous):

need hints? :D

OpenStudy (anonymous):

im thinking difference of two squares and difference of cubes but it doesnt make sense

OpenStudy (anonymous):

substitute" \[\LARGE a=\sqrt[3]{x} \quad \quad and\quad \quad b=1\] and you have: \[\LARGE a^3-b^3=(a-b)(a^2+ab+b^2 )\] so in your problem you have a-b so you need to multiply by a^2+ab+b^2 to get difference of cubes... (you'll have to do the same with numerator too. !) so they would cancel out!

OpenStudy (anonymous):

anyway, USE L'Hopital and tell me what do you get, so we can skip this boring limit :D

OpenStudy (anonymous):

USE L'Hopital...

OpenStudy (anonymous):

i thought you wont let me use it :P

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

lol how many times am i gonna lhopital this @_@

OpenStudy (anonymous):

not much... O_O show what you got: \[\LARGE \begin{array}{l}\mathop {\lim }\limits_{x \to 1} \left( {\frac{{\sqrt[4]{x} - 1}}{{\sqrt[3]{x} - 1}}} \right)' \\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{\frac{1}{{4\sqrt[4]{{{x^3}}}}}}}{{\frac{1}{{3\sqrt[3]{{{x^2}}}}}}}} \right)\\ = \left( {\frac{{\frac{1}{{4\sqrt[4]{{{1^3}}}}}}}{{\frac{1}{{3\sqrt[3]{{{1^2}}}}}}}} \right)\\ = \left( {\frac{{\frac{1}{4}}}{{\frac{1}{3}}}} \right) = \frac{3}{4}\end{array}\]

OpenStudy (anonymous):

:D ... YOUR TURN. ...

OpenStudy (anonymous):

YOU WIN! lol you made me feel incapable hahaha

OpenStudy (anonymous):

NO NO ... there's a thousand limits that we can get...Give me a real tough one, ... I always admited that my math sucks... but i want to learn ... don't feel bad for making a mistake which has not allowed you to get the answer. I had many of those, I'll have now I'm sure lol.... Go ahead your turn. ;)

OpenStudy (anonymous):

here's how it goes WITHOUT L'HOPITAL lol (now you'll know why I let you use L'Hopital)... I didn't want you to pass all through this. ;) (I wrote this on my programm while we were talking... ) isn't it Cute? \[\Large \begin{array}{l}\mathop {\lim }\limits_{x \to 1} \left( {\frac{{\sqrt[4]{x} - 1}}{{\sqrt[3]{x} - 1}}} \right)\\\\\sqrt[4]{x} = a\\1 = b\\{a^4} - {b^4} = \left( {a - b} \right)\left( {a + b} \right)\left( {{a^2} + {b^2}} \right)\\{a^4} - {b^4} = \left( {a - b} \right)\left( {{a^3} + a{b^2} + b{a^2} + {b^3}} \right)\\\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{a - b}}{{\sqrt[3]{x} - 1}} \cdot \frac{{{a^3} + a{b^2} + b{a^2} + {b^3}}}{{{a^3} + a{b^2} + b{a^2} + {b^3}}}} \right)\\\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{{a^4} - {b^4}}}{{\sqrt[3]{x} - 1\left( {{a^3} + a{b^2} + b{a^2} + {b^3}} \right)}}} \right)\\\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{{{\left( {\sqrt[4]{x}} \right)}^4} - {1^4}}}{{\sqrt[3]{x} - 1\left( {{{\left( {\sqrt[4]{x}} \right)}^3} + \sqrt[4]{x} + {{\left( {\sqrt[4]{x}} \right)}^2} + 1} \right)}}} \right)\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{x - 1}}{{\sqrt[3]{x} - 1\left( {{{\left( {\sqrt[4]{x}} \right)}^3} + \sqrt[4]{x} + {{\left( {\sqrt[4]{x}} \right)}^2} + 1} \right)}}} \right)\\\\now\\\sqrt[3]{x} = a\\1 = b\\{a^3} - {b^3} = \left( {a - b} \right)\left( {{a^2} + ab + {b^2}} \right)\\\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{x - 1}}{{a - b\left( {{{\left( {\sqrt[4]{x}} \right)}^3} + \sqrt[4]{x} + {{\left( {\sqrt[4]{x}} \right)}^2} + 1} \right)}}} \right) \cdot \left( {\frac{{\left( {{a^2} + ab + {b^2}} \right)}}{{\left( {{a^2} + ab + {b^2}} \right)}}} \right)\\\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{x - 1\left( {{a^2} + ab + {b^2}} \right)}}{{{a^3} - {b^3}\left( {{{\left( {\sqrt[4]{x}} \right)}^3} + \sqrt[4]{x} + {{\left( {\sqrt[4]{x}} \right)}^2} + 1} \right)}}} \right)\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{x - 1\left( {{{\left( {\sqrt[3]{x}} \right)}^2} + \sqrt[3]{x} + {1^2}} \right)}}{{{{\left( {\sqrt[3]{x}} \right)}^3} - {1^3}\left( {{{\left( {\sqrt[4]{x}} \right)}^3} + \sqrt[4]{x} + {{\left( {\sqrt[4]{x}} \right)}^2} + 1} \right)}}} \right)\\\\\mathop {\lim }\limits_{x \to 1} \left( {\frac{{x - 1\left( {{{\left( {\sqrt[3]{x}} \right)}^2} + \sqrt[3]{x} + {1^2}} \right)}}{{x - 1\left( {{{\left( {\sqrt[4]{x}} \right)}^3} + \sqrt[4]{x} + {{\left( {\sqrt[4]{x}} \right)}^2} + 1} \right)}}} \right)\\\\ = \left( {\frac{{\left( {{{\left( {\sqrt[3]{1}} \right)}^2} + \sqrt[3]{1} + {1^2}} \right)}}{{\left( {{{\left( {\sqrt[4]{1}} \right)}^3} + \sqrt[4]{1} + {{\left( {\sqrt[4]{1}} \right)}^2} + 1} \right)}}} \right)\\\\ = \left( {\frac{{1 + 1 + 1}}{{1 + 1 + 1 + 1}}} \right) = \frac{3}{4}\end{array}\]